Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
bersanetti, spinicelli, ruggi - 0:57 Tuesday 25 January 2022 (54593) Print this report
ISC Shift: lock robustness, BS/SR working point and test of new signals

This afternoon we continued the work on the robustness of the lock, testing the new signals which have been put online (see entries #54585#54590 and #54592). For this reason we did not test new features or additional loops, but we focused on a long lock and tried to understand the behaviour and reliability of such signals.

Here follows the summary of the activity; we had three short locks and a longer one, which lasted up to the end of the shift. The description of the four locks will follow; the other causes of unlocks during the shift were:

  • 17:07:07 UTC: unlock at the handoff of CARM to the MC;
  • 18:10:17 UTC and 18:14:29 UTC: unlock during the DRMI lock acquisition; to be checked offline, but they looked related to the ALS;
  • 18:46:01 UTC: unlock during the engagement of the SSFS.

About the four locks at CARM_NULL_1F:

  • the first lock at CARM_NULL_1F started at 16:18:40 UTC;
  • we started to optimize the working point with the usual methodology: offset on PR_TY, injection of DARM noise and tuning of the high frequency DARM pole, then the low frequency pole for the optical spring;
  • the dithering lines for the BS were turned on;
  • we wanted then to look the new signal LSC_NArm_B7_OS_I; we added the existing DARM_line_phi0 demodulation phase to the demodulation computation in LSC_Acl_Moni replacing the constant one;
  • we unlocked at 16:50:43 UTC when we restarted such process;

  • the second lock started at 17:18:02 UTC;
  • at 17:22:56 UTC we started to inject DARM noise and perform the DARM TF tuning; the high frequency pole was already in the proper place (DARM TF phase ~= 0.8 at 100 Hz);
  • we unlocked 17:38:57 UTC after a 0.2 urad movement on BS_TY destabilized the ITF;

  • the third lock started at 17:50:31 UTC;
  • we unlocked again at 18:01:53 UTC during the DARM TF tuning procedure, due to a strong ~13 Hz oscillation visible on all signals (Figure 1);

  • the fourth, and last, lock of the shift started at 19:02:20 UTC;
  • the overall trend of all figures of merit is in Figure 2;
  • this time we closed also the dithering loops, with smaller offsets than usual on X, while the Y DOFs were closed at zero; we had an increasing trend of the cavity powers during the lock;
  • we followed the usual procedure, but trying to use already the newly generated signals, and relying on the DARM noise injection only to cross-check the high frequency pole:
    • we used the LSC_DARM_HF_pole_mag signal to tune the high frequency pole; as reported in #54590 such signal represents the calibrated frequency once multiplied by 888; so we tried to tune it to be around 0.225 and cross-checking with the DARM noise injection wheter the DARM TF phase was approximately 0.8 at 100 Hz, by moving SR_TX and compensating with BS_TX for the dark fringe;
    • we used the LSC_OS_LF_mag_Hz signal to tune the optical spring pole down to 14 Hz, and looking at the behaviour of the new LSC_NArm_B7_OS_{I,Q} signals;
  • the entirety of the shift was devoted to keep the ITF locked and correct the alignment drifts of the BS and SR, in order to collect data and better characterize the new signals;
  • the demodulation phase of the new LSC_NArm_B7_OS_{I,Q} signals, i.e. LSC_DARM_line_phi0, was not optimized as it was not straightforward to remove any meaningful information from the _Q phase; we did a couple of attempts later towards the end of the shift, when we took some data with different demodulation phases; when we used 0.8 we found that the _Q signal was quite following the LSC_OS_LF_mag_Hz, so a change of pi/2 should have sufficed to bring the signal to _I; unfortunately for low pole frequencies the LSC_OS_LF_mag_Hz signal is not precise, and moreover we found at the end of the shift that the demodulation phase is not stored in the DAQ yet, so the tuning should be performed again; as a starting point one could set -0.77, which what we found during our tests (overall sign/pi to be checked); we added LSC_DARM_line_phi0 to the DAQ as a 1 Hz channel, but we did not restart LSC_Acl_Moni yet, as we did not want to unlock again as at the beginning of the shift; the process was restarted at the end of the activity;
  • during this last, long lock we witnessed four glitches on DARM (20:19:37 UTC, 20:43:47 UTC, 21:17:33 UTC and 21:37:23 UTC); they were relatively harmless as they did not unlock us, I wonder if they are of the same nature of the ones witnessed on Saturday when they were instead too much to handle for the LowNoise1 actuators.

At the end of the shift, to conclude the lock, we tested again the LowNoise1 configuration:

  • at 23:34:35 UTC we engaged the boost filters for both DARM and SRCL;
  • at 23:37:45 UTC we unlocked during the ramp of the LOWN_ENBL commands, due to a ~ 15 Hz oscillation on DARM (FIgure 3).

We left the arms locked on the IR.

Images attached to this report
Comments to this report:
flaminio - 2:18 Tuesday 25 January 2022 (54594) Print this report
I looked at the data collected during this shift.
The plot shows:
1. At the top the ITF signals: power on dark fringe, recycling cavity power, arm cavities power
2. In the middle the alignment of the north arm i.e. LSC_N*_T*_B7
3. At the bottom the alignment of the west arm i.e LSC_W*_T*_B8
There are four locks on the plot.
The first three are not very good locks. The alignment signals show some misalignment and the lock is lost.
During the last one, the crew managed to get the alignment signals at zero and the lock last long.
Images attached to this comment
Search Help